OpenStudy (anonymous):

Solve... Radical 3x/ Radical 2x + 3=2

OpenStudy (anonymous):

\[\sqrt{3x}\div \sqrt{2x+3}=2\]

OpenStudy (anonymous):

start with \[\frac{3x}{2x+3}=4\]

OpenStudy (anonymous):

What would be my first steps to solve this please?

OpenStudy (anonymous):

that is, square both sides

OpenStudy (anonymous):

Square both sides of the equation! Got it!

OpenStudy (anonymous):

then \[3x=4(2x+3)\]

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

don't forget to check at the end because you squared and that can mess things up. so check your solution make sure it is valid

OpenStudy (anonymous):

I got 12/-5, is that correct?

OpenStudy (anonymous):

yes but that means there is no solution

OpenStudy (anonymous):

why is there no solution?

OpenStudy (anonymous):

because you cannot take the square root of a negative number. so don't say "the answer is \[-\frac{12}{5}\] say there is no solution

OpenStudy (anonymous):

OK - Thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I have another one, I will post on the other side!